In a pure exchange economy with two goods, G and H, the two traders have Cobb-Douglas utility functions. Amos' utility is Ua= (Ga )^∝ (H)^(1-∝) and Elise’s is Ue= (Ge)^β (He)^(1-β). What are their marginal rates of substitution? Between them, Amos and Elise own 100 units of G and 50 units of H. Thus, If Amos has Ga and Ha, Elise has Ge =100-Ga and He =50-Ha


Solve for their contract curve. Show all your calculations clearly from your marginal utilities.



1
Expert's answer
2021-10-24T20:14:37-0400

Amos: rate of marginal substitution [MRSa = ["\alpha"/ (1-"\alpha") Ha/Ga

Elise: rate of marginal substitution [ MRSe = ["\beta" / (1- "\beta" ) He/Ge

the marginal rates of substitution are equal along the contract curves; MRSa= MRSe

equate the right-hand sides of the statements for MRSa and MRS_{e}MRSe

we finally make use of endowments information and some algebra to construct the contract curves quadratic formula respecting Amos goods to get the following equation;


("\beta"-"\alpha")G_{a}H_{a}+"\beta" ("\alpha"-1)50G_{a}+"\alpha" (1-"\beta")100H_{a}= 0("\beta"-"\alpha") GaHa+"\beta" (a-1)50Ga+"\alpha"(1-"\beta")100Ha=0

solve this way by substituting "\alpha" = "\beta"

therefore setting "\alpha" = "\beta"

contract curve becomes ("\beta" {2}-"\beta" )50G_{a} + ("\beta"-"\beta"{2})100H_{a}= 0("\beta"2-"\beta")50Ga+("\beta"-"\beta"2)100Ha=0

Divide by ("\beta"{2}-"\beta" ) ("\beta"2-"\beta" ) to get 50G_{a} -100H_{a} =50Ga-100Ha=0

we then use algebra to sum the equations. the straight line represents the contract curve t

Ga=2Ha